4R R 32H) Solve for Voltage Inductor using Impedances of Complet
Answer #1

o] given V. 32H XL = jaw - Vo=VsX Now Vo = V. X XL 4+ XL jaw - Vo=Vcx 4+j2w 2W/go Vo = us * 2 * 2 L tani paw] - vo z vs x Lao